OpenStudy (anonymous):

Simplify the following expression. sq rt -65 sq rt -70

OpenStudy (zepp):

\(\sqrt{-65}*\sqrt{-70}\)?

OpenStudy (anonymous):

there is no multiplication sign in the problem

OpenStudy (anonymous):

imaginary numbers

OpenStudy (jiteshmeghwal9):

am i wrong?

OpenStudy (zepp):

There's absolutely no imaginary numbers involved; \(\sqrt{-70*-65}\)

OpenStudy (zepp):

You'll get a positive number under the root, then you can simplify it.

OpenStudy (anonymous):

its -45 and -70 sorry....I wrote it wrong

OpenStudy (anonymous):

yeah there is, to simplify

OpenStudy (jiteshmeghwal9):

Yeah, zepp is right.

OpenStudy (zepp):

\(\sqrt{-70} \sqrt{-45}=\sqrt{-70*-45}=\sqrt{3150}\)

OpenStudy (jiteshmeghwal9):

plz tell me .Am I wrong?

OpenStudy (anonymous):

answer: -5 sq rt 195

OpenStudy (anonymous):

how did u get -45, clearly it is -65

OpenStudy (zepp):

http://puu.sh/BHro

OpenStudy (anonymous):

fine

OpenStudy (anonymous):

u win

OpenStudy (zepp):

You should get \(15\sqrt{14}\).

OpenStudy (anonymous):

I wrote it wrong...the question is -45 and -70

OpenStudy (anonymous):

actually ur wrong

OpenStudy (anonymous):

-15 sq rt 15

OpenStudy (lgbasallote):

@zepp remember!! \[\huge \sqrt{-a} \times \sqrt{-b} \implies i^2\sqrt{a \times b}\] ;)

OpenStudy (zepp):

Oh, I forgot a little minus? D:

OpenStudy (anonymous):

well, at least i knew there were imaginary numbers, but i^2 = -1

OpenStudy (zepp):

\(-15\sqrt{14}\) There :D

OpenStudy (zepp):

Looks alike anyway :P

OpenStudy (anonymous):

this question is confusing

OpenStudy (anonymous):

I agree!!

OpenStudy (anonymous):

zepp's answer is correct, case closed

OpenStudy (anonymous):

ok I understand up to 3150, then what do I do?

OpenStudy (anonymous):

its hard to explain... I just used my calculator

OpenStudy (anonymous):

I have to show my work lol

OpenStudy (anonymous):

it would be \[\sqrt{-70} i\]

OpenStudy (anonymous):

can u tell me how u get that because I am really confused

OpenStudy (anonymous):

ok use calculator. Do you have a scientific calculator?

OpenStudy (anonymous):

no and I have to show my work

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

does the problem says to write to hundredth or something?

OpenStudy (anonymous):

I multiplied -45 and -70 and got -3150, from here I am lost

OpenStudy (anonymous):

no, it just says simplify

OpenStudy (anonymous):

ok. since sqrt of 70 is not a integer, I leave it there and imaginary number is what you call sqrt of a negative number. To shows that, you do sqrt positive then i, indicating imaginary number. If this doesn't make sense to you, try asking others!:)

OpenStudy (anonymous):

ok thank you very much!!